Scientists want to place a 2600 kg satellite in orbit around Mars. They plan to have the satellite orbit at a speed of 2482 m/s in a perfectly circular orbit. Here is some information that may help solve this problem: mmars 4191 x 1023 kg rmars 3.397 x 106 m G 6.67428 x 10 N-m2/kg2 1) What radius should the satellite move at in its orbit? (Measured frrom t center of Mars.) he m Submit 2) What is the force of attraction between Mars and the satellite? N Submit 3) What is the acceleration of the satellite in orbit? m/s Submit 4) Which of the following quantities would change the radius the satellite needs to orbit at? he mass of the satellite he mass of the planet he speed of the satellite Submit 5) What should the speed of the orbit be, if we want the satellite to take 8 times longer to complete one full revolution of its orbit? m/s Submit
Answer #1


1)


In circular orbit


Fc = Fg

msatellite*v^2/d = G*msatellite*mmars/r^2


v =sqrt(G*mmars/r)

2482 = sqrt(6.67428*10^-11*6.4191*10^23/r)

r = 6.95*10^6 m


-------

2)

force of attraction F = G*mmars*msatellite/r^2

r = 6.95*10^6 m

F = 6.67428*10^-11*6.4191*10^23*2600/(6.95*10^6)^2

F = 2306 N


-----


3)

a = G*mmars/r^2 = 0.88 m/s^2

--------

4)

the speed of the satellite

5)

v = 2*pi*r/T


v2/v1 = T1/T2


T2 = 8*T1


v2 = v1/8 = 2482/8 = 310.25 m/s